OpenStudy (anonymous):

find the x- and y-intercepts: 5y-x=10

HanAkoSolo (jamierox4ev3r):

to find the x-intercept of this equation, you'd want to set the y to zero. Likewise, to find the y-intercept, you'd want to set the x to zero.

HanAkoSolo (jamierox4ev3r):

So for example, if you want to find your y-intercept, set the equation like this: 5y-0=10 ^ Here, we set the x to equal 0. Now, all you need to do is solve for y. Quite simple.

OpenStudy (anonymous):

so x-intercept is 2 and the y is -10?

HanAkoSolo (jamierox4ev3r):

other way around. the y-intercept is -10 and the x-intercept is 2. But yes, the numbers are correct. Nice job! :)

OpenStudy (anonymous):

thanks

HanAkoSolo (jamierox4ev3r):

No problem

OpenStudy (anonymous):

What about 2y+8x=-14?

HanAkoSolo (jamierox4ev3r):

Apply the same rule: \(\color{green}{\ast}\)to find the x-intercept, zero the y \(\color{green}{\ast}\)to find the y-intercept, zero the x Try it on your own, I'll check your answers

HanAkoSolo (jamierox4ev3r):

@tcorris do you need a hint?

OpenStudy (anonymous):

i have no idea

HanAkoSolo (jamierox4ev3r):

alright. So you do understand what I mean by "zero-ing" right? Basically, you just replace the offending variable with a zero To find the y-intercept, your equation becomes this: 2y+8(0)=-14 You can simplify that to become 2y=-14 To find the x-intercept, your equation becomes this: 2(0)+8x=-14 You can simplify that to become 8x=-14 Then once both equations are simplified down, all you need to do is solve for x (to find the x-intrecept) and solve for y (to find the y-intercept)

OpenStudy (anonymous):

I can't tell for the x but is the y -7?

HanAkoSolo (jamierox4ev3r):

the y-intercept is -7, that's correct

HanAkoSolo (jamierox4ev3r):

I've already given you the equation to find the x-intercept, which is 8x=-14 to solve this, divide both sides by 8 to isolate the x 8x=-14 /8 /8 ------- x=?? now all you have to do is divide -14 by 8. that is your x-intercept. Does that make sense?

OpenStudy (anonymous):

-1.75?

HanAkoSolo (jamierox4ev3r):

right! Just turn that into a fraction. which would be \(\Large\frac{-7}{4}\)

HanAkoSolo (jamierox4ev3r):

and that is your x-intercept. Do you understand now? All you have to do to find your x and y intercepts every single time is to zero the x and y variables in your equation. Hope this makes sense

OpenStudy (anonymous):

for 3x+4= y would the x be -1.3 and the y be 4?